OpenStudy (jungleman03):

Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three , solve for f(g(x)) when x = 2

OpenStudy (anonymous):

3.5+2\4 is 4.25

OpenStudy (anonymous):

there solved

OpenStudy (dayakar):

write equation of g(x)

OpenStudy (anonymous):

We do the same thing as your previous problem. substitute g(x) into every value of x in f(x). So.. \[f(g(x)) = (\frac{ x+4 }{ 3 })^2 + 2(\frac{x + 4}{3}) + 3.\] Now just plug 3 into x and evaluate it. \[f(g(3)) = (\frac{3+4}{3})^2 +2(\frac{3+4}{3}) + 3 = ?\]

OpenStudy (jungleman03):

x+4/3

OpenStudy (dayakar):

g(x) = (x+4)/3 if x=2 ind g(2) =

OpenStudy (dayakar):

find g(2) =

OpenStudy (dayakar):

r u doing

OpenStudy (jungleman03):

ya I'm trying but I'm still confused on how to set up the problem

OpenStudy (dayakar):

replace x with 2 in g(x) and simplify it

OpenStudy (dayakar):

show ur work